OpenStudy (anonymous):

medal(:

OpenStudy (anonymous):

For each function, y varies directly with x. Find each constant of variation. Then find the value of y when x= -0.3. 13. y= 2/3 when x= 0.2

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

\[y=kx\]and in this case \[\frac{2}{3}=0.2k\], so find \[k\](the constant of variation)then substitute it in the first equation i wrote to find \[y\]when \[x=-0.3\]

OpenStudy (anonymous):

oh okay so I would do 2/3=0.2(-0.3)?

OpenStudy (anonymous):

can you help me with math questions please? :) @pkjag

OpenStudy (anonymous):

No, you have to find a different \[y\], the first one has been given so that you can find \[k\]. First solve for \[k\] for now

OpenStudy (anonymous):

so instead of using the -0.3 I put in the 0.2 and solve?

OpenStudy (anonymous):

yes, just \[k\]for now

OpenStudy (anonymous):

okay im sorry im still a little confused /: can you give me step by step? Like the equation I start with and so on just so I know how to work other problems like this

OpenStudy (anonymous):

@Nnesha im confused do you think you could help me? /: its saying x=-0.3 but then in my question its giving me x=0.2

OpenStudy (anonymous):

we are working on the 2nd equation from my first reply

OpenStudy (anonymous):

just find\[k\] by dividing \[\frac{2}{3}\]with \[0.2\]

Nnesha (nnesha):

whats the question ???

OpenStudy (anonymous):

0.13?

OpenStudy (anonymous):

@Nnesha For each function, y varies directly with x. Find each constant of variation. Then find the value of y when x= -0.3. 13. y= 2/3 when x= 0.2

OpenStudy (anonymous):

see what throws me off is it says x=-0.3 and then the question says x=0.2? and no it doesn't have a table or anything that's why im confused /:

OpenStudy (anonymous):

2/3 divide by 0.2 isn't 0.13?

OpenStudy (anonymous):

@pkjag it is or isn't 0.13? cus when I did 2/3*0.2 that's what I got?

OpenStudy (anonymous):

it isn't do 2/0.6

OpenStudy (anonymous):

@pkjag I got 3.33? but how did you get 2/0.6?

OpenStudy (anonymous):

@nnesha you shouldn't be multiplying

OpenStudy (anonymous):

@Nnesha no I don't have multiple choice answers /: and I did 2/3*0.2 and I got 0.13

OpenStudy (anonymous):

yes now thats k the constant variation, use it to find y when x=-0.3

OpenStudy (anonymous):

use y=kx

OpenStudy (anonymous):

okay so y=3.33*-0.3?

OpenStudy (anonymous):

yes, but i don't understand part of your question: 'when x= -0.3. 13. y= 2/3 when x= 0.2' whats the 13?

OpenStudy (anonymous):

13 is just the question like 13. y= 2/3 when x=0.2 like this question that I may also need help with 15. y=4 when x=-3

OpenStudy (anonymous):

this:For each function, y varies directly with x. Find each constant of variation. Then find the value of y when x= -0.3. Is just saying what we need to follow for the questions I think .-. that's why im confused

OpenStudy (anonymous):

for all other questions like 13. start with finding the k, since you already have x and y, then use that k like you did in : 'okay so y=3.33*-0.3?'

Nnesha (nnesha):

 this is what i got so far